# InstructBioMol: Advancing Biomolecule Understanding and Design Following Human Instructions

### v1

@@ Results.
The performance of the generated protein enzymes across various evaluation metrics is presented in Figure 3d. InstructBioMol demonstrates the best performance in terms of similarity to ground truth, interaction capability with substrates, and exhibits superior generation validity. Specifically, InstructBioMol achieves improvements of 13.3 in ESP Score and 0.7 in Vina Score, indicating a stronger potential for substrate binding compared to baseline methods. Notably, InstructBioMol attains an ESP Score of 70.4, making it the only method to surpass the enzyme-substrate interaction threshold of 60.0 recommended by the ESP developer. This demonstrates that enzymes designed by InstructBioMol can bind their corresponding substrates with high affinity. Figure 4b in Supplementary Information further analyzes the top-1 ESP Score of the proteins generated for each substrate, revealing that InstructBioMol achieves the best performance on 66% of the substrates. Additionally, Supplementary Figure 4c presents the top-1 Vina Score on each substrate, with InstructBioMol attaining the best performance on 89% of the substrates. These findings suggest that InstructBioMol holds significant potential in generating highly efficient and specific protein enzymes, offering more effective solutions for fields such as biocatalysis.
Ablation analysis
Analysis of Training Strategy.
We first analyze the impact of different training strategies on model performance. Figure 4a presents a comparison of InstructBioMol with two of its variants: one variant retains continual pretraining and stage-2 instruction-tuning but removes stage-1 instruction-tuning (denoted as "w/ continual-pretraining, w/o stage-1 instruction-tuning"), while the other variant only retains stage-2 instruction-tuning but removes both continual pretraining and stage-1 instruction-tuning (denoted as "w/o continual-pretraining, w/o stage-1 instruction-tuning"). From these comparisons, we derive the following conclusions:
Firstly, in tasks related to molecule understanding and design (such as molecule captioning and description-based molecule generation), it is observed that the contribution of continual pretraining is relatively minor, whereas the significance of stage-1 instruction-tuning is more pronounced. In the molecule captioning task, removing stage-1 instruction-tuning results in an average performance drop of 7.3%, while further removal of continual pretraining leads to an additional drop of only 1.2%. In the molecule generation task, removing stage-1 instruction-tuning decreases the exact accuracy of generated molecules by 12.4%, and further removal of continual pretraining caused a subsequent drop of only 1.1%. We hypothesize that this is due to the critical role played by the molecular IUPAC name data used in stage-1 instruction-tuning, which effectively bridges and aligns molecular structure with natural language. Secondly, in description-based protein generation task, we find that contribution of continual pretraining is more significant, while the impact of stage-1 instruction-tuning is comparatively smaller. Specifically, removing stage-1 instruction-tuning leads to a decrease in the BLOSUM Substitution score by 0.187, and further removal of continual pretraining results in an additional drop of 0.326. This indicates that proteins, with their complex and intrinsic relationships such as sequence homology, benefit greatly from the evolutionary information captured by the broad protein sequence dataset used in the continual pretraining stage, thereby significantly enhancing the quality of the generated proteins.
Analysis of Multimodal Data Integration.
Next, we explore the impact of incorporating multimodal data on model performance. As shown in Figure 4b, we compare InstructBioMol and several of its variants on molecule captioning and protein function answering tasks. These variants include the removal of the 2D encoder (w/o 2D encoder) and 3D encoder (w/o 3D encoder) for molecular data, as well as the removal of the 1D encoder (w/o 1D encoder) and 3D encoder (w/o 3D encoder) for protein data. Additionally, we consider the removal of motif prompts (w/o motif) and the entire Motif-Guided Multimodal Feature Extraction Module (w/o multimodal). For comparison, the InstructBioMol variant used here excludes both the continual-pretraining stage and stage-1 instruction-tuning, serving as a basis for the ablation study on multimodal inputs.
The analysis results indicate that removing the Motif-Guided Multimodal Feature Extraction Module leads to a significant decline in model performance, with an average decrease of 4.7% and 7.1% on molecular and protein-related tasks, respectively. These findings underscore the importance of multimodal feature extraction in these tasks. Specifically, multimodal feature extraction can integrate information from different modalities, such as molecular structures and protein structures, thereby providing more comprehensive and accurate feature representations, which compensate for the limitations of single-modal features. Further analysis reveals that 3D data has a minimal impact on molecular-related tasks, with the removal of the 3D encoder resulting in only a 0.5% performance loss. Notwithstanding, for protein-related tasks, the influence of 3D data is more pronounced, with the removal of the 3D encoder causing a 4.2% decline in performance. This may be because the 3D information of molecules primarily describes the spatial distribution of atoms, overlooking detailed information about chemical bonds and functional groups. In contrast, the structure of proteins is fundamental to their functional performance, serving as a critical determinant of protein interactions and the formation of active sites. In conclusion, although different modalities contribute variably to different tasks, overall, multimodal feature extraction enhances the model’s performance in molecule and protein-related tasks. These results emphasize the importance of integrating multimodal data in biomolecular tasks.
@@ Methods
where $\mathbf{P}_{m}\in\mathbb{R}^{d}$ and $\mathbf{P}_{p}\in\mathbb{R}^{d}$ are motif prompt of molecule and protein, respectively. $\mathbf{M}_{m}\in\mathbb{R}^{N_{m}\times d}$ and $\mathbf{M}_{p}\in\mathbb{R}^{N_{p}\times d}$ are two learnable matrices. In detail, we obtain $\mathbf{T}_{m}$ by computing the Functional-Class FingerPrint (FCFP) [29] of a molecule with a radius of 2 and a length of 1024. The motifs of protein are collected from the UniProt [54] database, and details are described in Supplementary Information.
@@
The intrinsic complexity of biomolecules necessitates a modeling approach that is not only capable of capturing the detailed nuances of each modality but also adept at discerning the intricate interrelationships between them. Based on the extracted representations of a single modality, we propose to exploit the complementarities and redundancies between the various modalities to discover potentially more informative and robust representations than those obtained from any single-modality data. Herein, a Transformer [21] Encoder-Decoder architecture is employed for effective extraction and fusion of features from these heterogeneous sources. The reason for the choice of the Transformer architecture is that the inherent self-attention mechanism allows for the dynamic weighting of various parts of the input data, enabling to focus on the most relevant features across and within modalities. And the flexibility of the Encoder-Decoder facilitates the cross-modal integration of features, with the encoder capturing the salient features and the decoder synergizing those to construct a fused multimodal representation.
@@
By expanding the vocabulary and incorporating diverse multimodal features, we integrate molecules or proteins into textual formats, thereby augmenting the language model’s capacity to interpret biomolecules. Specifically, we concatenate these multimodal features $[\mathbf{z}_{1},\mathbf{z}_{2},...,\mathbf{z}_{(N_{q}+1)}]$ obtained from Eq. 5 with sequence-modality input, and label them with special tokens:
@@
$x_{multimodal\_mol}$ and $x_{multimodal\_prot}$ denote molecules and proteins, respectively. The inputs also encompass instructions $x_{insuction}$ in natural language, e.g., "What is the function of this protein", as well as description text $x_{text}$, e.g., "The molecule is a member of benzenes, a sulfone and a member of triazoles.". The composition of the inputs adapts based on the specific task. For instance, in generating molecular descriptions, the inputs include $x_{instruction}$ and $x_{multimodal}$, whereas in generating molecules from descriptions, the inputs consist of $x_{instruction}$ and $x_{text}$. The form of input corresponding to each task is detailed in Table 4.
Data type Entries Tokens Source
molecule 100M 4B PubChem
protein 49M 15B Uniref50
natural language 6M 8B PubMed, bioRxiv, ChemRxiv
Table 3: Statistics of continual pretraining dataset.
@@
We collect datasets on a hundred-million-scale, including a continual pretraining dataset (Table 3), and an instruction-tuning dataset (Table 4). The continual pretraining dataset comprises molecules and proteins in textual format, and natural language texts derived from scientific literature, enabling the model to develop a foundational adapting to biomolecular research. The instruction-tuning dataset contains various alignment pairs: molecule-natural language, protein-natural language, and molecule-protein (Figure 1d), achieving any-to-any alignment among molecules, proteins and natural language. Additionally, both molecular and protein data in the instruction-tuning dataset are multimodal, incorporating 2-D and 3-D structures of molecules and 3-D structures of proteins.
@@
For molecules, we collect 115 million entries from PubChem [56], filtering out those with atomic numbers exceeding 50, resulting in 100 million entries and 4 billion tokens after tokenization. For proteins, we use the Uniref50 dataset [57], which comprises 49 million entries and 15 billion tokens after tokenization. Recognizing the limitations of general language models trained on generalized corpora lacking biomolecular insights, we augment our data collection with literature specific to the research domain. Abstracts of scientific papers are collected across several sources, including PubMed [58], bioRxiv [59], and ChemRxiv [60], to enrich our training corpus. The incorporation of literature from these repositories enhances the domain-specific knowledge of InstructBioMol. This subset includes 6 million abstracts, resulting in 8 billion tokens.
@@
The dataset is sourced from two databases. The first is from PubChem, where we collect molecules with the IUPAC (International Union of Pure and Applied Chemistry) [61] name. This naming convention establishes a standardized nomenclature, fostering uniformity and clarity across the chemical community. 30 million molecules with IUPAC names are sampled from the filtered set in the previous step. The second source is ChEBI [22] data from ref. 14, comprising molecules alongside their descriptions. These descriptions encapsulate various facets of molecular structure, function, synthesis methodologies, etc.
@@
Data originate from two databases: SwissProt and TrEMBL [54]. These databases provide textual descriptions of proteins, covering four key aspects: name, family, location, and function. We utilize SwissProt data collected in ref. 19 and curate TrEMBL data from the UniProt [62] database. To ensure data quality and diversity, we filter TrEMBL dataset using UniRef50, resulting in approximately 25 million proteins. Additionally, for proteins described in at least three of four aspects, we consolidate these descriptions into a comprehensive summary using ChatGPT, creating data aligned from natural language to proteins.
@@
For molecule and protein pairs, we focus on two key applications: the discovery of drug-like molecules for specific target proteins and the design of enzyme proteins to catalyze specific substrates. Specifically, for generating molecules to specific target proteins, we use data from BindingDB [63] collected in ref. 64. BindingDB is a public database primarily focusing on the interactions between proteins, identified as potential targets, and small, drug-like molecular ligands. On the other hand, for designing an enzyme to catalyze a particular substrate, we draw upon data from the Rhea [65] database collected in ref. 20. Rhea is an expert-curated database of chemical reactions of biological interest, where enzyme-catalyzed reactions are curated from peer-reviewed literature.
@@
Despite the collection of a broad range of data, the quality of this data exhibits considerable variability across sources. For example, the ChEBI database offers a broader spectrum of molecular descriptions compared to the natural-language-like structure descriptions provided by IUPAC names in PubChem. Similarly, while data within the SwissProt undergo meticulous manual curation, entries in the TrEMBL do not benefit from such rigorous calibration. The tradeoff between the scale and quality of data poses significant challenges to model performance and generalizability. To address this issue, we adopt a two-stage instruction-tuning strategy designed to exploit the extensive data initially, then progressively direct the focus towards the insights offered by higher-quality datasets. Initially, in stage-1, the model is trained across all the available instructions. This stage leverages the diversity and volume of data to build a foundation on biomolecular alignment. Subsequently, in stage-2, the model undergoes further fine-tuning on a subset of higher-quality data. This approach harnesses both the expansive coverage of lower-quality data and the precision inherent in high-quality data, facilitating an efficient and effective utilization of the dataset. The scale and specific details of the different datasets used in the two stages are presented in Figure 1c and Table 4.
Sub-dataset Task type Scale Instruction stage-1 stage-2
Molecule - Natural Language Alignment
PubChem $(x_{instruction},x_{multimodal\_mol})\rightarrow y_{text}$ 30M Give the IUPAC name of the following molecule.
PubChem $(x_{instruction},x_{text})\rightarrow y_{mol}$ 30M Generate a molecule in SELFIES that fits the provided IUPAC name.
ChEBI $(x_{instruction},x_{multimodal\_mol})\rightarrow y_{text}$ 26K Provide a caption for the molecule below.
ChEBI $(x_{instruction},x_{text})\rightarrow y_{mol}$ 26K Generate a molecule in SELFIES that fits the provided description.
Protein - Natural Language Alignment
TrEMBL_Name $(x_{instruction},x_{multimodal\_protein})\rightarrow y_{text}$ 25M What is the official name of this protein?
TrEMBL_Family $(x_{instruction},x_{multimodal\_protein})\rightarrow y_{text}$ 5M What is the protein family that this protein belongs to?
TrEMBL_Locaction $(x_{instruction},x_{multimodal\_protein})\rightarrow y_{text}$ 2M What is the subcellular location of this protein?
TrEMBL_Function $(x_{instruction},x_{multimodal\_protein})\rightarrow y_{text}$ 1M What is the function of this protein?
TrEMBL_Description $(x_{instruction},x_{text})\rightarrow y_{prot}$ 2M Generate a protein matching the following description.
SwissProt_Name $(x_{instruction},x_{multimodal\_protein})\rightarrow y_{text}$ 455K What is the official name of this protein?
SwissProt_Family $(x_{instruction},x_{multimodal\_protein})\rightarrow y_{text}$ 370K What is the protein family that this protein belongs to?
SwissProt_Location $(x_{instruction},x_{multimodal\_protein})\rightarrow y_{text}$ 275K What is the subcellular location of this protein?
SwissProt_Function $(x_{instruction},x_{multimodal\_protein})\rightarrow y_{text}$ 411K What is the function of this protein?
SwissProt_Description $(x_{instruction},x_{text})\rightarrow y_{prot}$ 394K Generate a protein matching the following description.
Molecule – Protein Alignment
BindingDB $(x_{instruction},x_{multimodal\_prot})\rightarrow y_{mol}$ 335K Generate a drug molecule binding to the target protein.
Rhea $(x_{instruction},x_{multimodal\_mol})\rightarrow y_{prot}$ 190K Generate an enzyme that can catalyze for the given substrate.
Table 4: Statistics of instruction-tuning dataset.
@@
We use Pytorch [70] to implement the modal. The model is trained on 8 80G NVIDIA H800 GPUs. Additionally, we adopt the DeepSpeed ZeRO-1 [71] and BF16 for computational efficiency. The total number of training steps is 1.5 million. The steps for continual pretraining, stage-1 instruction-tuning and stage-2 instruction-tuning are 600,000, 500,000, and 400,000 respectively. We use the AdamW optimizer with $(\beta_{1},\beta_{2})$ set to $(0.9,0.95)$. We follow a linear learning rate schedule, warming up from 0 to maximum learning rate 1e-5 over the first 2,000 steps, and decaying the final learning rate down to 0. During training, all parameters are trainable except for the modality encoders $f_{m}^{2D}$, $f_{m}^{3D}$, $f_{p}^{1D}$, $f_{p}^{3D}$ in Equations 1 and 2, with the total trainable parameters being 6.8B. The ratio of different datasets sampled during training is controlled using hyper-parameters, which are detailed in Supplementary Table 1.
@@
To validate the model’s capability in molecule understanding and design, two tasks are performed: molecule captioning and description-based molecule generation. The dataset we use is consistent with ref. 14. To assess the model’s ability to understand and design proteins, we conduct experiments on protein property answering and description-based protein generation tasks. The data splits used in the experiments follow ref. 19 , with 3,000 samples selected as the test set for each task. Specifically, the protein properties include family, subcellular location, name, and function. For proteins containing at least three of these properties, we use ChatGPT to combine them into a complete description, which serves as the dataset for protein design based on the description. For the task of designing molecule drugs targeting proteins, we select 100 proteins as the test set. Similarly, for the task of designing enzymes based on substrates, we select 100 enzymes as the test set. Details of datasets are provided in the Supplementary Information Section 3.
@@
We leverage standard natural language generation metrics such as BLEU [23], ROUGE [24], and METEOR [25] to evaluate molecule captioning following ref. 14. These metrics measure how closely the generated captions match the reference captions.
@@
The following several types of metrics proposed in ref. 14 are used to evaluate the task of generating molecules from textual descriptions:
@@
Considering that both protein property answering and molecule captioning are natural language generation tasks, we adopt the evaluation metrics of the molecule captioning task to assess this task following ref. 14, 19.
@@
This metric is designed to measure the similarity between two protein sequences by calculating their percentage identity. It firstly counts the number of identical residues by comparing each corresponding residue in the reference protein $p_{ref}$ and the generated protein $p_{gen}$. Then it normalizes the number of identical residues by the sum of the lengths of both sequences. Formally,
@@
For drug discovery, we evaluate the generated molecules from three perspectives following ref. 37, 38:
@@
Following ref. 38. we use the Success Rate to assess the quality of the generated molecules by considering multiple factors, including binding affinity, drug-likeness, and synthetic accessibility. A molecule is successful if it meets specific thresholds for Vina Score, QED, and SA (Vina Score $<$ -8.18, QED $>$ 0.25, SA $>0.59$).
@@
In the task of enzyme design, since generations are protein sequences, we choose to use Indentity and Alignment metrics to assess the similarity between the generated and reference proteins. To assess the interaction between enzyme proteins and substrates, we employ two metrics: Vina Score and ESP Score [20]. Vina Score is used to quantify the strength of the interaction between the designed enzyme and its substrate. A lower value indicates a stronger interaction. Specifically, when calculating the Vina Score, we use DiffDock-L to obtain the complex structure and then use Qvina to obtain the corresponding score. ESP Score is another metric for evaluating enzyme-substrate interactions. This score ranges from 0 to 100, with higher scores indicating stronger interactions. We evaluate the model’s optimal performance by calculating the average top-1 Indentity, Alignment, Vina score, and ESP score for all substrate-specific designed proteins. Additionally, we use the Validity metric to evaluate the biological validity of the generated proteins.

### v2

@@ Results.
The performance of the generated protein enzymes across various evaluation metrics is presented in Figure 3d. InstructBioMol demonstrates the best performance in terms of similarity to ground truth, interaction capability with substrates, and exhibits superior generation validity. Specifically, InstructBioMol achieves improvements of 13.3% in ESP Score [24] and 0.7 in Vina Score (with $P$-values of $1.4\times 10^{-5}$ and $3.9\times 10^{-3}$), indicating a stronger potential for substrate binding compared to baseline methods. Notably, InstructBioMol attains an ESP Score of 70.4, making it the only method to surpass the enzyme-substrate interaction threshold of 60.0 recommended by the ESP developer. This demonstrates that enzymes designed by InstructBioMol can bind their corresponding substrates with high affinity. Supplementary Figure 1b further analyzes the top-1 ESP Score of the proteins generated for each substrate, revealing that InstructBioMol achieves the best performance on 66% of the substrates. Additionally, Supplementary Figure 1c presents the top-1 Vina Score on each substrate, with InstructBioMol attaining the best performance on 89% of the substrates. These findings suggest that InstructBioMol holds significant potential in generating highly efficient and specific protein enzymes, offering more effective solutions for fields such as biocatalysis.
Figure 4: Performance on description-based protein-molecule pairs generation. a, The generated proteins and molecules are evaluated from three perspectives: protein assessment (Alignment), molecule assessment (Vina Score, SA, QED), and overall assessment (Success Rate). b, -- Alignment × Vina Score across top-1, top-5, top-10, top-20, top-50, and all generated protein-molecule pairs for each description. A higher value indicates better quality of generated protein-molecule pairs.
InstructBioMol excels at designing protein and molecule simultaneously from descriptions
Experimental Setup.
As a unified Large Language Model, InstructBioMol has demonstrated exceptional alignment between natural language, molecules, and proteins. We further explore the advantages of this unified model by generating proteins and binding molecules simultaneously from textual descriptions. We split a training set and a test set, and the test set includes 100 target protein descriptions. In evaluation, for each description, 100 protein-molecule pairs are generated. The evaluation metrics are detailed in Methods. We adopt LoRA [38] fine-tuning on InstructBioMol for this task. Baselines consist of three groups. (1) GPT-3.5, which includes two variants: GPT-3.5 (5-shot, random) and GPT-3.5 (5-shot, similarity). (2) Fine-tuned Large Language Models. This includes Llama [7], Mol-Instructions-molecule, Mol-Instructions-protein, BioT5, and BioT5+. All of these models are fine-tuned on the training set. (3) Sequential ensemble approach: GPT-3.5 (5-shot, similarity) + DrugGPT, which combines the most competitive techniques for generating proteins from natural language and generating molecules from proteins. First, GPT-3.5 (5-shot, similarity) is used to generate proteins from descriptions, followed by the use of DrugGPT to generate molecules based on the proteins.
Results.
We present the experimental results in Figure 4, which demonstrate that InstructBioMol significantly outperforms the baseline methods. In protein generation, Alignment improves by 6.8%; in molecule generation, both Vina Score and SA achieve the best performance. Overall, the Success Rate reaches 48.3%, which is 15.2% higher than the baseline. Additionally, we analyze the top-n (-- Alignment $\times$ Vina Score) for each description, where a higher value indicates better quality of both generated proteins and molecules. We find that even in the top-20 results, our method still slightly outperforms the top-1 score of the strongest baseline, further proving the efficiency of InstructBioMol. We also observe that baselines based on fine-tuned Large Language Models do not perform as well. While Mol-Instructions-molecule, Mol-Instructions-protein, and InstructBioMol are all trained on Llama, their performances vary significantly. Specifically, Mol-Instructions-molecule, which is trained solely on text-molecule data, excels in molecule-related metrics like Vina Score, SA, and QED, but struggles with protein generation. In contrast, Mol-Instructions-protein, trained only on text-protein data, demonstrates strong protein generation capabilities with a high Alignment but performs poorly on molecule-related metrics. By comparison, InstructBioMol shows a marked performance advantage. This performance discrepancy clearly highlights that models specialized in either molecules or proteins struggle to meet the complex demands of tasks that require simultaneous integration of natural language, molecular, and protein data. Benefiting from achieving any-to-any alignment across natural language, molecules, and proteins, InstructBioMol can learn across different data types more effectively, capturing cross-modal semantic relationships.
@@ Methods
where $\mathbf{P}_{m}\in\mathbb{R}^{d}$ and $\mathbf{P}_{p}\in\mathbb{R}^{d}$ are motif prompt of molecule and protein, respectively. $\mathbf{M}_{m}\in\mathbb{R}^{N_{m}\times d}$ and $\mathbf{M}_{p}\in\mathbb{R}^{N_{p}\times d}$ are two learnable matrices. In detail, we obtain $\mathbf{T}_{m}$ by computing the Functional-Class FingerPrint (FCFP) [49] of a molecule with a radius of 2 and a length of 1024. The motifs of protein are collected from the UniProt [50] database, and details are described in Supplementary Information Section 1.
@@
A Transformer [25] Encoder-Decoder architecture is employed for effective extraction and fusion of features from biomolecular various modalities. The reason for the choice of the Transformer architecture is that the inherent self-attention mechanism allows for the dynamic weighting of various parts of the input data, enabling to focus on the most relevant features across and within modalities. The flexibility of the Encoder-Decoder facilitates the cross-modal integration of features, with the encoder capturing the salient features and the decoder synergizing those to construct a fused multimodal representation.
@@
For molecules and proteins, we concatenate multimodal features $\mathbf{Z}=[\mathbf{z}_{1},\mathbf{z}_{2},...,\mathbf{z}_{(N_{q}+1)}]$ obtained from Eq. 5 with sequence-modality input, and label them with special tokens:
@@
$x_{multimodal\_mol}$ and $x_{multimodal\_prot}$ denote molecules and proteins, respectively. In practice, we treat $\mathbf{Z}$ as an additional feature, concatenated with the text representation derived from tokenizing the input and passing it through the LLM embedding layer. The inputs also could include natural language instructions $x_{insuction}$, e.g., "What is the function of this protein", and description text $x_{text}$, e.g., "The molecule is a member of benzenes, a sulfone and a member of triazoles.". The specific composition of inputs depends on the task. For instance, generating molecular descriptions involves $x_{instruction}$ and $x_{multimodal}$, while generating molecules from descriptions uses $x_{instruction}$ and $x_{text}$. The detailed task-specific input forms are provided in Extended Data Table 2.
@@
We collect datasets on a hundred-million-scale, including a continual pretraining dataset (Extended Data Table 1), and an instruction-tuning dataset (Extended Data Table 2). The continual pretraining dataset comprises molecules and proteins in textual format, and natural language texts derived from scientific literature, enabling the model to develop a foundational adapting to biomolecular research. The instruction-tuning dataset contains various alignment pairs: molecule-natural language, protein-natural language, and molecule-protein (Figure 1d), achieving any-to-any alignment among molecules, proteins and natural language. Additionally, both molecular and protein data in the instruction-tuning dataset are multimodal, incorporating 2-D and 3-D structures of molecules and 3-D structures of proteins.
@@
For molecules, we collect 115 million entries from PubChem [52], filtering out those with atomic numbers exceeding 50, resulting in 100 million entries and 4 billion tokens after tokenization. For proteins, we use the Uniref50 dataset [53], which comprises 59 million entries and 15 billion tokens after tokenization. Recognizing the limitations of general language models trained on generalized corpora lacking biomolecular insights, we augment our data collection with literature specific to the research domain. Abstracts of scientific papers are collected across several sources, including PubMed [54], bioRxiv [55], and ChemRxiv [56], to enrich our training corpus. The incorporation of literature from these repositories enhances the domain-specific knowledge of InstructBioMol. This subset includes 6 million abstracts, resulting in 8 billion tokens.
@@
The dataset is sourced from two databases. The first is from PubChem, where we collect molecules with the IUPAC (International Union of Pure and Applied Chemistry) [57] name. This naming convention establishes a standardized nomenclature, fostering uniformity and clarity across the chemical community. 30 million molecules with IUPAC names are sampled from the filtered set in the previous step. The second source is ChEBI [26] data from ref. 15, comprising molecules alongside their descriptions. These descriptions encapsulate various facets of molecular structure, function, synthesis methodologies, etc.
@@
Data originate from two databases: SwissProt and TrEMBL [50]. These databases provide textual descriptions of proteins, covering four key aspects: name, family, location, and function. We utilize SwissProt data collected in ref. 20 and curate TrEMBL data from the UniProt [58] database. To ensure data quality and diversity, we filter TrEMBL dataset using UniRef50, resulting in approximately 25 million proteins. Additionally, for proteins described in at least three of four aspects, we consolidate these descriptions into a comprehensive summary using ChatGPT, creating data aligned from natural language to proteins.
@@
For molecule and protein pairs, we focus on two key applications: the discovery of drug-like molecules for specific target proteins and the design of enzyme proteins to catalyze specific substrates. Specifically, for generating molecules to specific target proteins, we use data from BindingDB [59] collected in ref. 60. BindingDB is a public database primarily focusing on the interactions between proteins, identified as potential targets, and small, drug-like molecular ligands. On the other hand, for designing an enzyme to catalyze a particular substrate, we draw upon data from the Rhea [61] database collected in ref. 24. Rhea is an expert-curated database of chemical reactions of biological interest, where enzyme-catalyzed reactions are curated from peer-reviewed literature.
@@
Despite the collection of a broad range of data, the quality of this data exhibits considerable variability across sources. For example, the ChEBI database offers a broader spectrum of molecular descriptions compared to the natural-language-like structure descriptions provided by IUPAC names in PubChem. Similarly, while data within the SwissProt undergo meticulous manual curation, entries in the TrEMBL do not benefit from such rigorous calibration. The tradeoff between the scale and quality of data poses significant challenges to model performance and generalizability. To address this issue, we adopt a two-stage instruction-tuning strategy designed to exploit the extensive data initially, then progressively direct the focus towards the insights offered by higher-quality datasets. Initially, in stage-1, the model is trained across all the available instructions. This stage leverages the diversity and volume of data to build a foundation on biomolecular alignment. Subsequently, in stage-2, the model undergoes further fine-tuning on a subset of higher-quality data. This approach harnesses both the expansive coverage of lower-quality data and the precision inherent in high-quality data, facilitating an efficient and effective utilization of the dataset. The scale and specific details of the different datasets used in the two stages are presented in Figure 1c and Extended Data Table 2.
@@
We use Pytorch [66] to implement the model. The model is trained on 8 80G NVIDIA H800 GPUs. Additionally, we adopt the DeepSpeed ZeRO-1 [67] and BF16 for computational efficiency. The total number of training steps is 1.5 million. The steps for continual pretraining, stage-1 instruction-tuning and stage-2 instruction-tuning are 600,000, 500,000, and 400,000 respectively. We use the AdamW optimizer with $(\beta_{1},\beta_{2})$ set to (0.9,0.95)(0.9,0.95)( 0.9 , 0.95 ). We follow a linear learning rate schedule, warming up from 0 to maximum learning rate 1e-5 over the first 2,000 steps, and decaying the final learning rate down to 0. During training, all parameters are trainable except for the modality encoders $f_{m}^{2D}$, $f_{m}^{3D}$, $f_{p}^{1D}$, $f_{p}^{3D}$ in Equations 1 and 2, with the total trainable parameters being 6.8B. The ratio of different datasets sampled during training is controlled using hyper-parameters, which are detailed in Supplementary Table 1.
@@
To validate the model’s capability in molecule understanding and design, two tasks are performed: molecule captioning and description-based molecule generation. The dataset we use is consistent with ref. 15. To assess the model’s ability to understand and design proteins, we conduct experiments on protein property answering and description-based protein generation tasks. The data splits used in the experiments follow ref. 20 , with 3,000 samples selected as the test set for each task. Specifically, the protein properties include family, subcellular location, name, and function. For proteins containing at least three of these properties, we use ChatGPT to combine them into a complete description, which serves as the dataset for protein design based on the description. For the task of designing molecule drugs targeting proteins, we select 100 proteins as the test set. Similarly, for the task of designing enzymes based on substrates, we select 100 enzymes as the test set. Details of datasets are provided in the Supplementary Information Section 3.
@@
We leverage standard natural language generation metrics such as BLEU [68], ROUGE [69], and METEOR [70] to evaluate molecule captioning following ref. 15. These metrics measure how closely the generated captions match the reference captions.
@@
The following several types of metrics proposed in ref. 15 are used to evaluate the task of generating molecules from textual descriptions:
@@
Considering that both protein property answering and molecule captioning are natural language generation tasks, we adopt the evaluation metrics of the molecule captioning task to assess this task following ref. 15, 20.
@@
This metric is designed to measure the similarity between two protein sequences by calculating their percentage identity. It first counts the number of identical residues by comparing each corresponding residue in the reference protein $p_{ref}$ and the generated protein $p_{gen}$. Then it normalizes the number of identical residues by the sum of the lengths of both sequences. Formally,
@@
For drug discovery, we evaluate the generated molecules from three perspectives following ref. 35, 78:
@@
Following ref. 78, we use the Success Rate to assess the quality of the generated molecules by considering multiple factors, including binding affinity, drug-likeness, and synthetic accessibility. A molecule is successful if it meets specific thresholds for Vina Score, QED, and SA (Vina Score $<$ -8.18, QED $>$ 0.25, SA >0.59>0.59> 0.59).
@@
In the task of enzyme design, since generations are protein sequences, we choose to use Identity and Alignment metrics to assess the similarity between the generated and reference proteins. To assess the interaction between enzyme proteins and substrates, we employ two metrics: Vina Score and ESP Score [24]. Vina Score is used to quantify the strength of the interaction between the designed enzyme and its substrate. A lower value indicates a stronger interaction. Specifically, when calculating the Vina Score, we use DiffDock-L to obtain the complex structure and then use Qvina to obtain the corresponding score. ESP Score is another metric for evaluating enzyme-substrate interactions. This score ranges from 0 to 100, with higher scores indicating stronger interactions. We evaluate the model’s optimal performance by calculating the average top-1 Identity, Alignment, Vina score, and ESP score for all substrate-specific designed proteins. Additionally, we use the Validity metric to evaluate the biological validity of the generated proteins.
Evaluation Metrics for Designing Protein and Molecule from Descriptions
The evaluation metrics are organized into three groups: (1) Protein evaluation: This focuses on calculating the Alignment between the generated proteins and the ground truth. (2) Molecule evaluation: This group evaluates the generated molecules using the Vina Score to assess the binding affinity between the molecules and proteins, as well as assesses molecular properties including QED and SA. (3) Overall evaluation: This group combines both generated proteins and molecules. Specifically, we use the Success Rate to assess the proportion of successful generations, considering factors such as Alignment, Vina Score, QED, and SA. A generation is considered successful if it meets the following specific thresholds: Alignment $>$ 30%, Vina Score $>$ -8.18, QED $>$ 0.25, and SA $>$ 0.59.
Data availability
Dataset used in this study is available in Zenodo at https://doi.org/10.5281/zenodo.15303508 (ref. 83).
Code availability
The source code of this study is available in GitHub at https://github.com/HICAI-ZJU/InstructBioMol (ref. 84).
Acknowledgements
This work is funded by NSFCU23B2055 (H.C.), NSFC2302433 (Q.Z.), NSFCU23A20496 (Q.Z.), the Fundamental Research Funds for the Central Universities (226-2023-00138, H.C.), Zhejiang Provincial “Jianbing” “Lingyan” Research and Development Program of China (2025C01097, K.D. and Q.Z.), Zhejiang Provincial Natural Science Foundation of China (LQ24F020007, Q.Z.) and Hangzhou West Lake Pearl Project Leading Innovative Youth Team Project (TD2023017, K.D.).
Author Contributions Statement
X.Z., K.D., Q.Z., and H.C. conceived the study. X.Z. developed the method, implemented the code, and conducted the experiments. T.L. participated in benchmarking some baseline models. X.Z., Y.J., X.L., and Z.X. contributed to dataset collection. K.D., Z.W., M.Q., K.F., J.W., Q.Z., and H.C. provided critical suggestions on methodology and experiments. All authors wrote the paper, reviewed it, and approved the final paper.
Competing Interests Statement
The authors declare no competing interests.